# Central simple algebras, Milnor $K$-theory and homogeneous spaces over complete discretely valued fields of dimension 2

### v1

@@ 4.   Symbols and Milnor $K$-theory
We prove now injectivity of the restriction using corestriction, which corresponds to the norm morphism $N_{K^{\prime}/K}:\mathrm{k}^{\mathrm{M}}_{2}(K^{\prime})\to\mathrm{k}^{\mathrm{M}}_{2}(K)$. Since $(d,p)=1$, we know that the composite ${\rm{Cores\,}}\circ{\rm{Res\,}}$ is a bijection on every $u_{2}^{i}(K)$. In particular, we know that it defines a bijection on the quotient $u_{2}^{i}(K)/u_{2}^{i+1}(K)$. It will suffice then to prove that this map factors through $u_{2}^{di}(K^{\prime})/u_{2}^{di+1}(K^{\prime})$, which amounts to proving that $N_{K^{\prime}/K}$ sends $u_{2}^{di+1}(K^{\prime})$ to $u_{2}^{i+1}(K)$. Note that this is evident for $i=\epsilon_{K}$ since $di=\epsilon_{K^{\prime}}$ and hence $u_{2}^{di+1}(K^{\prime})=0$, so we assume henceforth that $i<\epsilon_{K}$.
Keeping our choice of $\pi^{\prime}=\sqrt[d]{\pi}$, we know by isomorphisms (7) and (8) that any element in $u_{2}^{di+1}(K^{\prime})$ can be written as a sum of symbols of the form
  • $\{1-{\pi^{\prime}}^{j}x_{j},y_{j}\}$ (if $(j,p)=1$ and $\omega=-\pi^{\prime j}$);
  • $\{1-{\pi^{\prime}}^{j}x_{j},\pi^{\prime}\}$ (if $p|j$ and $\varpi=-\pi^{\prime j/p}$);
  • and an element in $u_{2}^{di+d}(K^{\prime})$;
where $di+1\leq j\leq di+d-1$ and $x_{j},y_{j}\in\mathcal{O}_{K}$. And since we already proved that the restriction is surjective on the quotients $u_{2}^{j}(K^{\prime})/u_{2}^{j+1}(K^{\prime})$ for any $j$, we know that $u_{2}^{di+d}(K^{\prime})$ is sent to $u_{2}^{i+1}(K)$ by the corestriction, so we may focus on the remaining symbols.
Let us compute the norm of an element of the form $1-{\pi^{\prime}}^{j}x$ with $x\in\mathcal{O}_{K}$. Write $e=(j,d)$, $d=ef$ and $j=eg$. Then the conjugates of ${\pi^{\prime}}^{j}$ are precisely $\zeta_{f}^{k}{\pi^{\prime}}^{j}$ for $10\leq k\leq f-$ 1. Thus, we get
$N_{K^{\prime}/K}(1-{\pi^{\prime}}^{j}x)=\left(\prod_{k=0}^{f-1}(1-\zeta_{f}^{k}{\pi^{\prime}}^{j}x)\right)^{e}=(1-{\pi^{\prime}}^{fj}x^{f})^{e}=(1-{\pi^{\prime}}^{feg}x^{f})^{e}=(1-{\pi}^{g}x^{f})^{e}.$
Consider now a symbol of the form $\{1-{\pi^{\prime}}^{j}x,\pi^{\prime}\}$ with $x\in\mathcal{O}_{K}$ and $p|j$. Then, if we fix $m$ such that $md\equiv 1\pmod{p}$, we have
$\{1-{\pi^{\prime}}^{j}x,\pi^{\prime}\}=\{1-{\pi^{\prime}}^{j}x,{\pi^{\prime}}^{md}\}=\{1-{\pi^{\prime}}^{j}x,\pi^{m}\}=m\{1-{\pi^{\prime}}^{j}x,\pi\}.$
@@
$N_{K^{\prime}/K}(\{1-{\pi^{\prime}}^{j}x,\pi^{\prime}\})=N_{K^{\prime}/K}(m\{1-{\pi^{\prime}}^{j}x,\pi\})=m\{N_{K^{\prime}/K}(1-{\pi^{\prime}}^{j}x),\pi\}\\ =m\{(1-{\pi}^{g}x^{f})^{e},\pi\}=me\{1-{\pi}^{g}x^{f},\pi\}.$ (12)
On the other hand, consider a symbol in $u_{2}^{j}(K^{\prime})$ of the form $\{1-{\pi^{\prime}}^{j}x,y\}$ with $x,y\in\mathcal{O}_{K}$ and $(j,p)=1$. Then, taking corestriction we have
$N_{K^{\prime}/K}(\{1-{\pi^{\prime}}^{j}x,y\})=\{N_{K^{\prime}/K}(1-{\pi^{\prime}}^{j}x),y\}=\{(1-{\pi}^{g}x^{f})^{e},y\}=e\{1-{\pi}^{g}x^{f},y\}.$ (13)
@@
Having proved that (10) is an isomorphism, fix an arbitrary uniformizer $\pi\in K$. Write $\pi={\pi^{\prime}}^{d}u_{1}$ for some uniformizer $\pi^{\prime}\in\mathcal{O}_{K^{\prime}}$ and some $u_{1}\in\mathcal{O}_{K^{\prime}}^{\times}$. Then:
@@
  • We put $\rho_{0}:={\rm{Res\,}}^{-1}_{K^{\prime}/K}\circ m_{d}\circ\rho_{0}^{\prime}$, where $m_{d}$ denotes multiplication by $d$ and where we use the uniformizer $\pi^{\prime}$ to define $\rho_{0}^{\prime}$. Then we see that, for a lift $y\in\mathcal{O}_{K}$ of $\bar{y}\in\bar{K}$ we have
    $\rho_{0}(\bar{y})={\rm{Res\,}}^{-1}_{K^{\prime}/K}(d\{y,\pi^{\prime}\})={\rm{Res\,}}^{-1}_{K^{\prime}/K}(\{y,{\pi^{\prime}}^{d}\})=\{y,\pi\},$
@@
    where the last equality uses the fact that $\pi={\pi^{\prime}}^{d}u_{1}$ and $\{y,u_{1}\}\in u_{2}^{1}(K)$ by isomorphism (1) and the fact that $\mathrm{k}^{\mathrm{M}}_{2}(\bar{K})=0$.
@@
  • For $0<i<\epsilon_{K}$ with $(i,p)=1$ and $\omega\in K$ such that $v_{K}(\omega)=i$, we put $\rho_{i}:={\rm{Res\,}}^{-1}_{K^{\prime}/K}\circ\rho_{di}^{\prime}$, where we use $\omega$ as the element with $v_{K^{\prime}}(\omega)=di$. Then we see that, for lifts $x,y\in\mathcal{O}_{K}$ of $\bar{x},\bar{y}\in\bar{K}$ we have
    $\rho_{i}\left(\bar{x}\frac{d\bar{y}}{\bar{y}}\right)={\rm{Res\,}}^{-1}_{K^{\prime}/K}(\{1+\omega x,y\})=\{1+\omega x,y\}.$
@@
  • For $0<i<\epsilon_{K}$ with $p|i$ and $\varpi\in K$ with $v_{K}(\varpi)=i/p$, we put $\rho_{i}:={\rm{Res\,}}^{-1}_{K^{\prime}/K}\circ m_{d}\circ\rho_{di}^{\prime}$, where we use $\pi^{\prime}$ as the uniformizer and $\varpi$ as the element with $v_{K^{\prime}}(\varpi)=di/p$. Then we see that, for a lift $x\in\mathcal{O}_{K}$ of $\bar{x}\in\bar{K}$ we have
    $\rho_{i}\left(\bar{x}\pmod{\bar{K}^{p}}\right)={\rm{Res\,}}^{-1}_{K^{\prime}/K}(d\{1+\varpi^{p}x,\pi^{\prime}\})={\rm{Res\,}}^{-1}_{K^{\prime}/K}(\{1+\varpi^{p}x,{\pi^{\prime}}^{d}\})=\{1+\varpi^{p}x,\pi\},$
@@
    where the last equality uses the fact that $\pi={\pi^{\prime}}^{d}u_{1}$ and $\{1+\varpi^{p}x,u_{1}\}\in u_{2}^{i+1}(K)$ by isomorphism (3) and the fact that $\Omega_{\bar{K}}^{1}/\mathrm{Ker}(d)\subseteq\Omega_{\bar{K}}^{2}=0$.
  • For $i=\epsilon_{K}$, we put $\rho_{i}:={\rm{Res\,}}^{-1}_{K^{\prime}/K}\circ m_{d}\circ\rho_{di}^{\prime}$, where we use $\pi^{\prime}$ as the uniformizer to define $\rho^{\prime}_{di}$. Then we see that, for a lift $x\in\mathcal{O}_{K}$ of $\bar{x}\in\bar{K}$ we have
    $\rho_{i}\left([\bar{x}]\right)={\rm{Res\,}}^{-1}_{K^{\prime}/K}(d\{1+(\zeta_{p}-1)^{p}x,\pi^{\prime}\})={\rm{Res\,}}^{-1}_{K^{\prime}/K}(\{1+(\zeta_{p}-1)^{p}x,{\pi^{\prime}}^{d}\})\\ ={\rm{Res\,}}^{-1}_{K^{\prime}/K}(\{1+\pi^{\epsilon_{K}}ux,{\pi^{\prime}}^{d}\})=\{1+\pi^{\epsilon_{K}}ux,\pi\},$
    where $u$ is defined as in (11) and the last equality uses the fact that $\pi={\pi^{\prime}}^{d}u_{1}$ and $\{1+\pi^{\epsilon_{K}}ux,u_{1}\}\in u_{2}^{\epsilon_{K}+1}(K)$ by isomorphism (4) and the fact that $H_{p}^{2}(\bar{K})=0$.
@@
Now that we have extended Kato’s isomorphisms, we can follow the same proof structure as for [LPS, Prop. 1.8], which we used in the case where $K$ was assumed to contain all $p$-th roots of unity. Indeed, to do so, one may settle the following three “absorption lemmas” that are analoguous to Lemmas 1.5, 1.6 and 1.7 of loc. cit. and that will be helpful to inductively reduce every Milnor $\mathrm{K}$-theory class to a single symbol:
@@
If $i=\epsilon_{K}$, then isomorphism (9) already tells us that our class is a symbol since $H^{1}_{p}(\bar{K})$ is a quotient of $\Omega_{\bar{K}}^{0}=\bar{K}$ and hence every element in $H^{1}_{p}(\bar{K})$ has the form $[\bar{x}]$ for some $\bar{x}\in\bar{K}$.
If $i<\epsilon_{K}$, fix a uniformizer $\pi$ and put $\omega:=-\pi^{i}$ if $(i,p)=1$ and $\varpi:=-\pi^{i/p}$ if $p|i$. Then we may apply one of the isomorphisms (6), (7) or (8) and see that $\alpha=a+b$, where $a$ is a symbol in $u_{2}^{i}(K)\smallsetminus u_{2}^{i+1}(K)$ (given by the corresponding isomorphism), and $b\in u_{2}^{i+1}(K)$. We may apply then inductively one of the Lemmas 4.4, 4.5 or 4.6, so that $\alpha=a+c$, with $a$ a symbol as above and $c\in u_{2}^{\lceil\epsilon_{K}\rceil}(K)$. If $\epsilon_{K}\not\in{\mathbb{Z}}$, then we are done since we know that $u_{2}^{\lceil\epsilon_{K}\rceil}(K)=0$ in this case by Proposition 4.3. Otherwise, we have two cases:
If $p|i$ (including $i=0$), then $a=\{y,\pi\}$ for some $y\in\mathcal{O}_{K}$ and some uniformizer $\pi\in K$. On the other hand, isomorphism (9) tells us that $c=\{1-\pi^{\epsilon_{K}}ux,\pi\}$ for some $x,u\in\mathcal{O}_{K}$. We have then
$\alpha=a+c=\{y(1-\pi^{\epsilon_{K}}ux),\pi\}.$
@@
Putting then $\pi^{\prime}:=\pi x^{m}y$ and applying isomorphism (9) to this uniformizer, we see that $c=\{1-\pi^{\epsilon_{K}}uz,\pi^{\prime}\}$ for some $z,u\in\mathcal{O}_{K}$. We have then
$\alpha=a+c=\{(1-\pi^{i}x)(1-\pi^{\epsilon_{K}}uz),\pi^{\prime}\},$

### v2

@@ 4.   Symbols and Milnor $K$-theory
We prove now injectivity of the restriction using corestriction, which corresponds to the norm map $N_{M/K}:\mathrm{k}^{\mathrm{M}}_{2}(M)\to\mathrm{k}^{\mathrm{M}}_{2}(K)$. Since $(d,p)=1$, we know that the composite ${\rm{Cores\,}}\circ{\rm{Res\,}}$ is a bijection on every $u_{2}^{i}(K)$. In particular, we know that it defines a bijection on the quotient $u_{2}^{i}(K)/u_{2}^{i+1}(K)$ and thus $N_{M/K}$ will correspond to the inverse of ${\rm{Res\,}}_{M/K}$ up to multiplication by $d$ once we prove that it is bijective. Since ${\rm{Cores\,}}\circ{\rm{Res\,}}$ is a bijection and ${\rm{Res\,}}$ is surjective on $u_{2}^{di}(M)/u_{2}^{di+1}(M)$, it will suffice to prove that ${\rm{Cores\,}}\circ{\rm{Res\,}}$ factors through $u_{2}^{di}(M)/u_{2}^{di+1}(M)$, which amounts to proving that $N_{M/K}$ sends $u_{2}^{di+1}(M)$ to $u_{2}^{i+1}(K)$. Note that this is evident for $i=\epsilon_{K}$ since $di=\epsilon_{M}$ and hence $u_{2}^{di+1}(M)=0$, so we assume henceforth that $i<\epsilon_{K}$.
Keeping our choice of $\pi_{M}=\sqrt[d]{\pi}$, we know by isomorphisms (6) and (7) that any element in $u_{2}^{di+1}(M)$ can be written as a sum of symbols of the form
  • $\{1-\pi_{M}^{j}x_{j},y_{j}\}$ (if $(j,p)=1$ and $\omega=-\pi_{M}^{j}$);
  • $\{1-\pi_{M}^{j}x_{j},\pi_{M}\}$ (if $p|j$ and $\varpi=-\pi_{M}^{j/p}$);
  • and an element in $u_{2}^{di+d}(M)$;
where $di+1\leq j\leq di+d-1$ and $x_{j},y_{j}\in\mathcal{O}_{K}$. And since we already proved that the restriction is surjective on the quotients $u_{2}^{j}(M)/u_{2}^{j+1}(M)$ for any $j$, we know that $u_{2}^{di+d}(M)$ is sent to $u_{2}^{i+1}(K)$ by the corestriction, so we may focus on the remaining symbols.
Let us compute the norm of an element of the form $1-\pi_{M}^{j}x$ with $x\in\mathcal{O}_{K}$. Write $e=(j,d)$, $d=ef$ and $j=eg$. Then the conjugates of $\pi_{M}^{j}$ are precisely $\zeta_{f}^{k}\pi_{M}^{j}$ for $10\leq k\leq f-$ 1. Thus, we get
$N_{M/K}(1-\pi_{M}^{j}x)=\left(\prod_{k=0}^{f-1}(1-\zeta_{f}^{k}\pi_{M}^{j}x)\right)^{e}=(1-\pi_{M}^{fj}x^{f})^{e}=(1-\pi_{M}^{feg}x^{f})^{e}=(1-{\pi}^{g}x^{f})^{e}.$
Consider now a symbol of the form $\{1-\pi_{M}^{j}x,\pi_{M}\}$ with $x\in\mathcal{O}_{K}$ and $p|j$. Then, if we fix $m$ such that $md\equiv 1\pmod{p}$, we have
$\{1-\pi_{M}^{j}x,\pi_{M}\}=\{1-\pi_{M}^{j}x,\pi_{M}^{md}\}=\{1-\pi_{M}^{j}x,\pi^{m}\}=m\{1-\pi_{M}^{j}x,\pi\}.$
@@
$N_{M/K}(\{1-\pi_{M}^{j}x,\pi_{M}\})=N_{M/K}(m\{1-\pi_{M}^{j}x,\pi\})=m\{N_{M/K}(1-\pi_{M}^{j}x),\pi\}\\ =m\{(1-{\pi}^{g}x^{f})^{e},\pi\}=me\{1-{\pi}^{g}x^{f},\pi\}.$ (10)
On the other hand, consider a symbol in $u_{2}^{j}(M)$ of the form $\{1-\pi_{M}^{j}x,y\}$ with $x,y\in\mathcal{O}_{K}$ and $(j,p)=1$. Then, taking corestriction we have
$N_{M/K}(\{1-\pi_{M}^{j}x,y\})=\{N_{M/K}(1-\pi_{M}^{j}x),y\}=\{(1-{\pi}^{g}x^{f})^{e},y\}=e\{1-{\pi}^{g}x^{f},y\}.$ (11)
@@
Having proved that (9) is an isomorphism, fix an arbitrary uniformizer $\pi\in K$. Write $\pi=\pi_{M}^{d}u_{1}$ for some uniformizer $\pi_{M}\in\mathcal{O}_{M}$ and some $u_{1}\in\mathcal{O}_{M}^{\times}$. Then
@@
  • We put $\rho_{0}:={\rm{Res\,}}^{-1}_{M/K}\circ m_{d}\circ\rho_{0}^{M}$, where $m_{d}$ denotes multiplication by $d$ and where we use the uniformizer $\pi_{M}$ to define $\rho_{0}^{M}$. Then we see that, for a lift $y\in\mathcal{O}_{K}$ of $\bar{y}\in\bar{K}$ we have
    $\rho_{0}(\bar{y})={\rm{Res\,}}^{-1}_{M/K}(d\{y,\pi_{M}\})={\rm{Res\,}}^{-1}_{M/K}(\{y,\pi_{M}^{d}\})=\{y,\pi\},$
@@
    where the last equality uses the fact that $\pi=\pi_{M}^{d}u_{1}$ and $\{y,u_{1}\}\in u_{2}^{1}(K)$ by isomorphism (1) and the fact that $\mathrm{k}^{\mathrm{M}}_{2}(\bar{K})=0$.
@@
  • For $0<i<\epsilon_{K}$ with $(i,p)=1$ and $\omega\in K$ such that $v_{K}(\omega)=i$, we put $\rho_{i}:={\rm{Res\,}}^{-1}_{M/K}\circ\rho_{di}^{M}$, where we use $\omega$ as the element with $v_{M}(\omega)=di$. Then we see that, for lifts $x,y\in\mathcal{O}_{K}$ of $\bar{x},\bar{y}\in\bar{K}$ we have
    $\rho_{i}\left(\bar{x}\frac{d\bar{y}}{\bar{y}}\right)={\rm{Res\,}}^{-1}_{M/K}(\{1+\omega x,y\})=\{1+\omega x,y\}.$
@@
  • For $0<i<\epsilon_{K}$ with $p|i$ and $\varpi\in K$ with $v_{K}(\varpi)=i/p$, we put $\rho_{i}:={\rm{Res\,}}^{-1}_{M/K}\circ m_{d}\circ\rho_{di}^{M}$, where we use $\pi_{M}$ as the uniformizer and $\varpi$ as the element with $v_{M}(\varpi)=di/p$. Then we see that, for a lift $x\in\mathcal{O}_{K}$ of $\bar{x}\in\bar{K}$ we have
    $\rho_{i}\left(\bar{x}\pmod{\bar{K}^{p}}\right)={\rm{Res\,}}^{-1}_{M/K}(d\{1+\varpi^{p}x,\pi_{M}\})={\rm{Res\,}}^{-1}_{M/K}(\{1+\varpi^{p}x,\pi_{M}^{d}\})=\{1+\varpi^{p}x,\pi\},$
@@
    where the last equality uses the fact that $\pi=\pi_{M}^{d}u_{1}$ and $\{1+\varpi^{p}x,u_{1}\}\in u_{2}^{i+1}(K)$ by isomorphism (3) and the fact that $\Omega_{\bar{K}}^{1}/\mathrm{Ker}(d)\subseteq\Omega_{\bar{K}}^{2}=0$.
  • For $i=\epsilon_{K}$, we consider the composite $\theta_{i}:=(\rho_{di}^{M})^{-1}\circ{\rm{Res\,}}_{M/K}$, where we use $\pi_{M}$ as the uniformizer. The morphism $\theta_{i}$ is injective, and we want to compute its image. To do so, observe that the group $u_{2}^{i}(K)$ is generated by symbols of the form $\{1+\omega x,y\}$ and $\{1+\omega x,\pi\}$ with $x,y\in\mathcal{O}_{K}^{\times}$. Since ${\rm{Res\,}}_{M/K}$ is injective and every symbol of the form $\{1+\omega x,y\}$ with $x,y\in\mathcal{O}_{M}^{\times}$ is trivial in $u_{2}^{i}(M)$, we deduce that $u_{2}^{i}(K)$ is generated by symbols of the form $\{1+\omega x,\pi\}$ with $x,y\in\mathcal{O}_{K}^{\times}$. Hence the image of $\theta_{i}$ is generated by the elements of the form
    $\theta_{i}(\{1+\omega x,\pi\})=\theta_{i}(d\{1+(1+\zeta_{p})^{p}u^{-1}x,\pi_{M}\})=d\bar{u}^{-1}\bar{x},$
    with $x,y\in\mathcal{O}_{K}^{\times}$. Here, the first equality uses the fact that $\pi=\pi_{M}^{d}u_{1}$ and $\{1+\omega x,u_{1}\}\in u_{2}^{\epsilon_{M}+1}(M)$ by isomorphism (4) and the fact that $H^{2}_{p}(\bar{M})=0$. We therefore have an isomorphism
    $\eta_{i}:\bar{K}/(\bar{K}\cap\overline{u}\Phi(\bar{M}))$ $\to{\rm{Im\,}}(\theta_{i})\subseteq\bar{M}/\Phi(\bar{M})=H_{p}^{1}(\bar{M})$
    $\overline{x}$ $\mapsto d\overline{u}^{-1}\overline{x},$
    and we set $\rho_{i}:=\theta_{i}^{-1}\circ\eta_{i}$, which sends,$[\bar{x}]$ to $\{1+\omega x,\pi\}$.
This concludes the proof of the case where $L=K$.
We move on to the general case. Since $M=L(\zeta_{p})$ and $M/L$ is totally ramified, we have $\bar{L}=\bar{M}$ and the isomorphisms (5)–(8) at the level of $L$, which we denote by $\rho_{i}^{L}$.
Note that, since $L/K$ is unramified, we have ${\rm{Res\,}}_{L/K}(u_{2}^{i}(K))\subseteq u_{2}^{i}(L)$. Thus there is a natural map
${\rm{Res\,}}_{L/K}:u_{2}^{i}(K)/u_{2}^{i+1}(K)\to u_{2}^{i}(L)/u_{2}^{i+1}(L),$
which is injective for every $0\leq i\leq\epsilon_{K}$ by the classical restriction-corestriction argument.
We define then $\rho_{i}$ as follows. Fix an arbitrary uniformizer $\pi\in K$ and note that it corresponds to a uniformizer of $L$ as well. Then
  • For $i=0$, we define $\rho_{0}:={\rm{Res\,}}_{L/K}^{-1}\circ\rho_{0}^{L}\circ\iota$, where $\iota:\bar{K}\to\bar{L}$ denotes the inclusion and where we use the uniformizer $\pi$ to define $\rho_{0}^{L}$. This is well defined since, for a lift $y\in\mathcal{O}_{K}$ of $\bar{y}\in\bar{K}$ we have
    $\rho_{0}^{L}\circ\iota(\bar{y})=\{y,\pi\},$
    which is clearly in the image of ${\rm{Res\,}}_{L/K}$ and coincides with the morphism in (5). It is injective since $\iota$ is injective. For the surjectivity, observe that the group $\mathrm{k}^{\mathrm{M}}_{2}(K)/u^{1}_{2}(K)$ is spanned by elements of the form $\{y,z\}$ or $\{y,\pi\}$ with $y$ and $z$ in $\mathcal{O}_{K}^{\times}$. But the restriction to $\mathrm{k}^{\mathrm{M}}_{2}(L)/u^{1}_{2}(L)$ of an element of the form $\{y,z\}$ with $y,z\in\mathcal{O}_{K}^{\times}$ is trivial, and hence, by injectivity of the restriction, the symbol $\{y,z\}$ itself is trivial in $\mathrm{k}^{\mathrm{M}}_{2}(K)/u^{1}_{2}(K)$. We deduce that $\mathrm{k}^{\mathrm{M}}_{2}(K)/u^{1}_{2}(K)$ is spanned by elements of the form $\{y,\pi\}$ with $y\in\mathcal{O}_{K}^{\times}$ and such an element is obviously in the image of $\rho_{0}$.
  • For $0<i<\epsilon_{K}$ with $(i,p)=1$ and $\omega\in K$ such that $v_{K}(\omega)=i$, we define $\rho_{i}:={\rm{Res\,}}_{L/K}^{-1}\circ\rho_{i}^{L}\circ\iota_{1}$, where $\iota_{1}:\Omega^{1}_{\bar{K}}\to\Omega^{1}_{\bar{L}}$ denotes the natural map induced by the inclusion $\iota$ and where we use $\omega$ as the element with $v_{L}(\omega)=i$. This is well defined since, for lifts $x,y\in\mathcal{O}_{K}$ of $\bar{x},\bar{y}\in\bar{K}$ we have
    $\rho_{i}^{L}\circ\iota_{1}\left(\bar{x}\frac{d\bar{y}}{\bar{y}}\right)=\{1+\omega x,y\}.$
    which is clearly in the image of ${\rm{Res\,}}_{L/K}$ and coincides with the morphism in (6). It is injective since $\iota_{1}$ is injective. For the surjectivity, observe that the group $u^{i}_{2}(K)/u^{i+1}_{2}(K)$ is spanned by elements of the form $\{1+\omega x,y\}$ and $\{1+\omega x,\pi\}$ with $x$ and $y$ in $\mathcal{O}_{K}^{\times}$. But for each element of the form $\{1+\omega x,\pi\}$ with $x\in\mathcal{O}_{K}^{\times}$, if we take $r$ such that $ri+1\equiv 0\mod p$, we can rewrite the symbol as
    $\{1+\omega x,\pi\}=\{1+\omega x,(-\omega x)^{r}\pi\}=\{1+\omega x,u\pi^{ri+1}\}=\{1+\omega x,u\},$
    for some unit $u\in\mathcal{O}_{K}^{\times}$. Hence $u^{i}_{2}(K)/u^{i+1}_{2}(K)$ is spanned by elements of the form $\{1+\omega x,y\}$ with $x$ and $y$ in $\mathcal{O}_{K}^{\times}$ and such an element is obviously in the image of $\rho_{i}$.
  • For $0<i<\epsilon_{K}$ with $p|i$ and $\varpi\in K$ with $v_{K}(\varpi)=i/p$, we put $\rho_{i}:={\rm{Res\,}}^{-1}_{L/K}\circ\rho_{i}^{L}\circ\iota_{0}$, where $\iota_{0}:\bar{K}/\bar{K}^{p}=\Omega_{\bar{K}}^{0}/\mathrm{Ker}(d)\to\Omega_{\bar{L}}^{0}/\mathrm{Ker}(d)=\bar{L}/\bar{L}^{p}$ denotes the natural map induced by the inclusion $\iota$ and where we use $\pi$ as the uniformizer and $\varpi$ as the element with $v_{M}(\varpi)=i/p$. This is well-defined since, for a lift $x\in\mathcal{O}_{K}$ of $\bar{x}\in\bar{K}$ we have
    $\rho_{i}^{L}\circ\iota_{0}\left(\bar{x}\pmod{\bar{K}^{p}}\right)=\rho_{i}^{L}\left(\bar{x}\pmod{\bar{L}^{p}}\right)=\{1+\varpi^{p}x,\pi\},$
    which is clearly in the image of ${\rm{Res\,}}_{L/K}$ and coincides with the morphism in (7). Since $\bar{L}/\bar{K}$ is separable, the map $\iota_{0}$ is injective and hence so is $\rho_{i}$. For the surjectivity, observe that the group $u^{i}_{2}(K)/u^{i+1}_{2}(K)$ is spanned by elements of the form $\{1+\varpi^{p}x,y\}$ and $\{1+\varpi^{p}x,\pi\}$ with $x$ and $y$ in $\mathcal{O}_{K}^{\times}$. But each element of the form $\{1+\varpi^{p}x,y\}$ with $x,y\in\mathcal{O}_{K}^{\times}$ is trivial since it becomes trivial in $u^{i}_{2}(L)/u^{i+1}_{2}(L)$ and the restriction map $u^{i}_{2}(K)/u^{i+1}_{2}(K)\rightarrow u^{i}_{2}(L)/u^{i+1}_{2}(L)$ is injective. We deduce that $u^{i}_{2}(K)/u^{i+1}_{2}(K)$ is spanned by elements of the form $\{1+\varpi^{p}x,\pi\}$ with $x\in\mathcal{O}_{K}^{\times}$ and such an element is obviously in the image of $\rho_{i}$.
  • For $i=\epsilon_{K}$, we consider the composite $\theta_{i}:=(\rho_{i}^{L})^{-1}\circ{\rm{Res\,}}_{L/K}$. The morphism $\theta_{i}$ is injective, and we want to compute its image. To do so, observe that the group $u_{2}^{i}(K)$ is generated by symbols of the form $\{1+\omega x,y\}$ and $\{1+\omega x,\pi\}$ with $x,y\in\mathcal{O}_{K}^{\times}$. Since ${\rm{Res\,}}_{L/K}$ is injective and every symbol of the form $\{1+\omega x,y\}$ with $x,y\in\mathcal{O}_{L}^{\times}$ is trivial in $u_{2}^{i}(L)$, we deduce that $u_{2}^{i}(K)$ is generated by symbols of the form $\{1+\omega x,\pi\}$ with $x,y\in\mathcal{O}_{K}^{\times}$. Hence the image of $\theta_{i}$ is generated by the elements of the form $\theta_{i}(\{1+\omega x,\pi\})=\bar{x}$ with $x\in\mathcal{O}_{K}^{\times}$. We therefore have an isomorphism
    $\eta_{i}:\bar{K}/(\bar{K}\cap\overline{u}\Phi(\bar{M}))$ $\to{\rm{Im\,}}(\theta_{i})\subseteq\bar{L}/(\bar{L}\cap\overline{u}\Phi(\bar{M}))=\bar{M}/\Phi(\bar{M})$
    $\overline{x}$ $\mapsto\overline{x},$
    and we set $\rho_{i}:=\theta_{i}^{-1}\circ\eta_{i}$, which sends $[\bar{x}]$ to $\{1+\omega x,\pi\}$.
This concludes the proof in the general case. ∎
@@
Now that we have extended Kato’s isomorphisms, we can follow the same proof structure as for [LPS, Prop. 1.8], which we used in the case where $K$ was assumed to contain all $p$-th roots of unity. Indeed, to do so, one may settle the following three “absorption lemmas” that are analoguous to Lemmas 1.5, 1.6 and 1.7 of loc. cit. and that will be helpful to inductively reduce every Milnor $\mathrm{K}$-theory class to a single symbol.
@@
If $i=\epsilon_{K}$, then isomorphism (8) already tells us that our class is a symbol.
If $i<\epsilon_{K}$, fix a uniformizer $\pi$ and put $\omega:=-\pi^{i}$ if $(i,p)=1$ and $\varpi:=-\pi^{i/p}$ if $p|i$. Then we may apply one of the isomorphisms (5), (6) or (7) and see that $\alpha=a+b$, where $a$ is a symbol in $u_{2}^{i}(K)\smallsetminus u_{2}^{i+1}(K)$ (given by the corresponding isomorphism), and $b\in u_{2}^{i+1}(K)$. We may apply then inductively one of the Lemmas 4.5, 4.6 or 4.7, so that $\alpha=a+c$, with $a$ a symbol as above and $c\in u_{2}^{\lceil\epsilon_{K}\rceil}(K)$. If $\epsilon_{K}\not\in{\mathbb{Z}}$, then we are done since we know that $u_{2}^{\lceil\epsilon_{K}\rceil}(K)=0$ in this case by Proposition 4.3. Otherwise, we have two cases:
If $p|i$ (including $i=0$), then $a=\{y,\pi\}$ for some $y\in\mathcal{O}_{K}$ and some uniformizer $\pi\in K$. On the other hand, isomorphism (8) tells us that $c=\{1-\omega x,\pi\}$ for some $x,\omega\in\mathcal{O}_{K}$. We have then
$\alpha=a+c=\{y(1-\omega ux),\pi\}.$
@@
Putting then $\pi^{\prime}:=\pi x^{m}y$ and applying isomorphism (8) to this uniformizer, we see that $c=\{1-\omega z,\pi^{\prime}\}$ for some $z,\omega\in\mathcal{O}_{K}$. We have then
$\alpha=a+c=\{(1-\pi^{i}x)(1-\omega z),\pi^{\prime}\},$